### v3

# Feeding the Crave: How People with Eating Disorders Get Trapped in the Perpetual Cycle of Digital Food Content

Abstract.
Recent studies have examined how digital food content impacts viewers’ dietary health. A few have found that individuals with eating disorders are particularly sensitive to digital food content, such as eating and cooking videos, which contribute to disordered eating behaviors. However, there is a lack of comprehensive studies that investigate how these individuals interact with various digital food content. To fill this gap, we conducted two rounds of studies (N=23 and 22, respectively) with individuals with eating disorders to understand their motivations and practices of consuming digital food content. Our study reveals that participants anticipate positive effects from food media to overcome their condition, but in practice, it often exacerbates their disorder. We also discovered that many participants experienced a cycle of quitting and returning to digital food content consumption. Based on these findings, we articulate design implications for digital food content and multimedia platforms to support vulnerable individuals.
copyright: none
1. Introduction
Caution: This paper discusses eating disorders and includes media that could potentially be a trigger to those dealing with eating disorders. Please use discretion when reading and disseminating this paper.
Digital media’s exponential growth has made food content ubiquitous and diversified. On social media, the popularity of food videos (e.g., eating and cooking broadcasts) and images has significantly increased (Kang et al., 2020; Mejova et al., 2016; De Solier, 2005; Adema, 2000). On TikTok, the hashtag #TikTokFood has been viewed over 43 billion times (Kitchens, 2021), and Instagram has over 310 million posts tagged with #foodporn (Instagram, 2024). The online food market, especially food delivery services often showcasing visually enticing images, has also seen tremendous growth, more than threefold, since 2017, rising as a global market worth more than \$150 billion (Insights, 2023; Inc., 2021).
A growing body of research has examined the impact of increased exposure to various digital food content on eating habits and dietary health (Petit et al., 2016; Kinard, 2016; Ofli et al., 2017; Vermeir and Roose, 2020; Pope et al., 2015; Choi et al., 2024; Wu et al., 2024). The trend of watching and posting food images and videos with catchy hashtags has been associated with inadequate eating habits (Yan and Corino, 2019; Qutteina et al., 2019), unhealthy food choices (Peng and JEMMOTT III, 2018; Basso et al., 2018), and overeating (Bodenlos and Wormuth, 2013; Halford et al., 2004) through various visual and social factors (Seal et al., 2022). Browsing and ordering food using online shopping and food delivery services is also known to involve unhealthy food choices (Yeo et al., 2017), unhealthy eating (Eu and Sameeha, 2021), and obesity (Stephens et al., 2020). Moreover, many digital food content features take-away food (Kang et al., 2020), and combining eating content with online food delivery services could lead to unhealthy lifestyles.
Recent studies have highlighted that the negative impacts of digital food content on dietary health are more intense in individuals with eating disorders (EDs), who are particularly sensitive to such content (Jin, 2018; Legenbauer et al., 2004) and engage with it significantly more than the general population (Bachner-Melman et al., 2018). When coupled with frequent unconscious exposure to digital food content, digital food content could exacerbate ED symptoms, such as binge eating and obsessive food thoughts (Choi et al., 2024; Kinkel-Ram et al., 2022). Despite these findings, there remains a significant gap in understanding how individuals with EDs engage with digital food content and how such engagement ultimately contributes to their disordered eating in their daily lives. In this regard, we aim to shed light on the experience of people with disordered eating behaviors with digital food content by answering the following question: How do people with EDs perceive and interact with digital food content?
To answer this question, we conducted two rounds of studies with individuals with eating disorders: one round of exploratory study (N=23) and another round of in-depth study (N=22). The exploratory study assesses how many individuals are concerned with associations of digital food content and disordered eating, and identifies which digital food content types are associated with disordered eating. Notably, we refrained from bringing up digital food content, allowing participants to broach the subject voluntarily. Sixteen participants out of twenty-three voluntarily linked digital food content to their ED symptoms. Our analysis identified two representative digital food content positively associated with ED symptoms: food media (e.g., images and videos) and food delivery apps.
To deepen our understanding of ED individuals’ motivation and practices of consuming each type of digital food content, we conducted in-depth interviews with 22 individuals (including two from our exploratory study) with EDs. Our main finding highlights a significant disparity between the expectations of individuals when using digital food content (e.g., vicarious satisfaction) and the actual outcomes they experience (e.g., a binge-eating stimulus). Amidst this disparity, many individuals have experienced a recurring cycle of quitting and returning to digital food content consumption behavior. In addition, we identified a unique subcategory of food media that some utilized as a self-care strategy: videos created by people with EDs (e.g., binge-eating broadcasts)
Our finding emphasizes the significant impact of digital food exposure on individuals with EDs, often resulting in developing or exacerbating disordered eating behaviors. The increasing presence of virtual food potentially triggering unhealthy eating habits fails to consider the growing occurrence of eating disorders, which is a concern for relevant organizations and policymakers (Taillie et al., 2019; Times, 2023; Rilloraza, 2024). Understanding the significance of visually and auditorily presented food in today’s context is crucial for guiding populations toward more appropriate food behaviors and choices. From our findings, we propose design implications for intervention to promote healthy interactions with digital food content.
The main contributions of this paper are summarized as follows.
  • We identified two representative kinds of digital food content, food media and food delivery apps, associated with eating disorders in everyday life.
  • Through an in-depth study, we revealed the perceptions, motivations, impacts, and patterns of consuming digital food content among individuals with eating disorders.
  • Finally, we provided design implications to improve multimedia platforms for vulnerable populations and offered insights for future directions in digital food content curation.
  • Our study complements prior research on digital food content by focusing on individuals with eating disorders, providing unique insights into their digital content consumption patterns and implications for HCI designers.
2. Related Work
2.1. Eating Disorders and Digital Content Use
The intersection of digital content use and eating disorders has been a focal point in the HCI community (Feuston et al., 2020; Pater et al., 2021, 2019). Studies have identified that extensive engagement with digital media, including pro-eating disorder content and imagery showcasing idealized body images, correlates with the onset and exacerbation of eating disorders (Chancellor et al., 2017; Gak et al., 2022). Research has also underpinned the critical role of digital environments, which often emphasize thinness and perfection as desirable body standards, in shaping individuals’ perceptions of body image and eating behaviors (Aparicio-Martinez et al., 2019). In addition, such social media environments encouraged individuals to constantly compare individuals with influencers and celebrities, leading to negative body image and low self-esteem, driving individuals to engage in disordered eating behaviors to meet unrealistic ideals (Pedalino and Camerini, 2022). Individuals vulnerable to body dissatisfaction and disordered eating behaviors are particularly susceptible to these media messages, as they fuel the internalization of harmful beauty ideals (MurrayStuart et al., 2018). The consistent engagement in such content contributed to the persistence and intensification of eating disorder symptoms.
The interplay between digital content and eating disorders has underscored the need for a comprehensive investigation that considers the multifaceted nature of the impact of digital content on eating disorders (Chancellor et al., 2016a, b). While a substantial body of research has explored the link between various digital content consumption and eating disorders, there is a notable gap in our understanding of the specific association between food-related content, which is closely connected to dietary health (Petit et al., 2016; Kinard, 2016; Ofli et al., 2017; Vermeir and Roose, 2020), and disordered eating behaviors.
2.2. An Overarching Presence of Digital Food Content
Digital food content comprises diverse multimedia materials that revolve around the food theme. This encompasses various content types, including captivating recipe videos, tantalizing food imagery, informative cooking tutorials, and engaging food blogs. Moreover, the proliferation of food delivery apps that present visually captivating food pictures has expanded digital food content, providing users with seamless access to various culinary offerings through their digital devices (Maimaiti et al., 2018).
In today’s digital age, the prevalence of digital food content is palpable across various online platforms (Kang et al., 2020; Mejova et al., 2016; Weber et al., 2021; Kitchens, 2021). Individuals constantly engage in various food-related content, from social media to video-sharing platforms, dedicated food websites, and mobile applications. This content not only caters to culinary enthusiasts seeking inspiration and knowledge (De Solier, 2005; Wang et al., 2022) but also to a broader audience drawn by the visual allure and cultural narratives associated with food (Spence et al., 2016; Adema, 2000; Peng and JEMMOTT III, 2018). As a result, the pervasive presence of digital food content has transformed how people engage with and perceive food in their daily lives (Yan and Corino, 2019).
This omnipresence of digital food content underscores the need to explore its impact on eating behaviors, especially among individuals with eating disorders. The subsequent section delves into the intricate dynamics of digital food content’s influence on individuals’ eating behaviors, shedding light on its constructive and detrimental effects.
2.3. Double-sided Impact of Digital Food Content on Eating Behaviors
Digital food content exerts a complex influence on individuals’ eating behaviors, presenting a dual-sided impact characterized by both constructive and detrimental facets (Strand and Gustafsson, 2020). While many individuals engage with digital food content for inspiration, recipe ideas, and cultural connections to food (De Solier, 2005), the comprehensive impact of such media on individuals with eating disorders, who are much more sensitive to digital food content than the general population (Schienle et al., 2009; Lee et al., 2017), has received limited attention. The relationship between individuals’ consumption of digital food content, including the engagement with eating broadcasts (Mukbang), and the potential escalation of disordered eating behaviors has garnered recognition in prior research, primarily focusing on the general population (von Ash et al., 2023; Anari and Eghtesadi, 2023). However, these investigations have yielded varying insights and have yet to explicitly target individuals with eating disorders.
Certain instances of digital food content consumption have been associated with fostering positive eating habits and healthier dietary choices (Drivas et al., 2024). Informative cooking tutorials, recipe demonstrations, and nutrition education videos can empower individuals to experiment with wholesome meal options and develop improved culinary skills (Wang et al., 2022; Ventura et al., 2021). These platforms potentially serve as educational tools, equipping users with knowledge to make informed decisions about their food intake. Such exposure can lead to a sense of agency in managing dietary preferences and restrictions, encouraging healthier relationships with food (Donnar, 2017). Furthermore, for certain individuals, engaging with digital food media provides a sense of presence that helps alleviate their loneliness (Anjani et al., 2020). However, it is essential to acknowledge that these potential positive effects are not uniformly experienced, and their effectiveness can depend on individual and personal contexts (Strand and Gustafsson, 2020; Seal et al., 2022).
Conversely, a growing body of research has highlighted the potentially detrimental impact of digital food content, especially when exacerbating unhealthy eating habits (Pope et al., 2015; Spence et al., 2016). The omnipresence of visually appealing, calorie-laden food imagery in digital content can escalate disordered eating behaviors (Bodenlos and Wormuth, 2013; Ventura et al., 2021). Studies have pointed out that exposure to images and videos of excessive food consumption, particularly those featuring eating challenges or binge eating, can inadvertently trigger or reinforce unhealthy eating patterns (Spence et al., 2019). Such content inadvertently glorifies or normalizes overconsumption (Kircaburun et al., 2021c). Additionally, the visual appeal of enticing food displays and the sensory experience of hearing eating sounds may inadvertently intensify cravings and compulsive eating tendencies (Spence et al., 2016; Anjani et al., 2020).
In addition, food delivery apps, despite their convenience, have been implicated in negatively influencing eating habits (Portingale et al., 2023). These platforms often feature visually appealing images of indulgent and calorie-rich foods strategically designed to entice users (Kapoor and Vij, 2018). Such visual appeals can lead to impulsive and unplanned consumption online (Liu et al., 2013). Moreover, frequent use of food delivery apps can normalize excessive consumption and encourage a sedentary lifestyle, further exacerbating the risk of weight gain and related health issues (Stephens et al., 2020; Maimaiti et al., 2018). The targeted marketing tactics employed by these apps might amplify the temptation to order food impulsively, making it challenging for individuals, especially those susceptible to disordered eating tendencies, to maintain a balanced and mindful approach to their eating habits.
The current research landscape lacks a profound understanding of how individuals with eating disorders engage with digital food content—a realm encompassing spontaneous interactions and passive influences. Existing studies are confined to quantitative survey-based methods, such as examining the correlation between the presence of eating disorders and the frequency of viewing eating broadcast (Kircaburun et al., 2021c, a; von Ash et al., 2023; Kim et al., 2021; Anari and Eghtesadi, 2023; Bachner-Melman et al., 2018). While these studies are meaningful in establishing links between eating disorders and digital food content, they leave the deeper psychological and behavioral dynamics largely unexplored, creating a significant gap in our understanding. These studies suggested further research into the various aspects of digital food content and media types that potentially promote disordered eating (Kircaburun et al., 2021a; Kim et al., 2021). The goal of our research is to shed light on these intricate dynamics and unravel the profound impact these interactions have on individuals’ attitudes, behaviors, and overall well-being. We aim to provide insights into the mechanisms underpinning the sway of digital food content on this population, paving the way for effective and tailored interventions and support mechanisms on digital food content platforms.
3. Exploratory Study
Before delving deeply into the investigation of how digital food content affects individuals with EDs, we conducted an IRB-approved exploratory study to broadly assess the extent to which they connect with digital food content in relation to their ED symptoms.
3.1. Participants
We recruited 23 participants (aged 18-43, mean=26.2 years; 19 females, 1 male, 1 non-binary, 2 chose not to disclose) through an advertisement post on an online community for people with EDs under permission of the moderators in South Korea  (cop., 2018). The participant gender composition is predominantly female, reflecting universal ED demographics (87.4% of identified South Korean binge eating disorder patients being female (Service, 2023), and threefold higher lifetime prevalence of binge eating disorder and bulimia nervosa in females in the US (Hudson et al., 2007; Galmiche et al., 2019)). We provide an overview of participant demographics, ED information, and types of digital food content mentioned by each participant in our Supplementary Materials. Our study intentionally included individuals who were either formally diagnosed or self-diagnosed. By doing so, we aimed to encompass a population that may not have sought professional treatment, providing insights into the often overlooked experiences of these individuals within a clinical context.
3.2. Study Protocol
The study involved an online survey to investigate how people with EDs associate digital food content with their ED symptoms. Additionally, we conducted remote semi-structured interviews for a deeper understanding. We employed the Eating Disorder Examination Questionnaire 6.0 (EDE-Q) to assess if our participants represented the eating disorder population (Fairburn and Beglin, 2008). [^1] Higher scores on the EDE-Q indicate problematic eating behaviors and attitudes. The EDE-Q scores of our study’s interviewees (4.56±0.56) were generally higher than those of individuals diagnosed with ED (4.02±1.28) (Aardoom et al., 2012). To address the social stigma associated with EDs (Puhl and Suh, 2015), participants were given the option to opt out of any part of the study, which consists of a survey, an interview, and the EDE-Q, they find uncomfortable.
In surveys, we intentionally refrained from introducing digital food content to accurately estimate how many participants consider exposure to digital food content to affect their symptoms. Instead, we asked about their computing device usage associated with disordered eating compared with normal meals. This approach allowed us to identify unique device usages, including consumption of digital food content, associated with their disordered eating. Survey questions included are: Is there any digital device usage associated with the time before/during/after eating disorder episodes compared with the time before/during/after usual meals? If so, which device(s) do you use and for what purpose?
Each Zoom interview, lasting approximately thirty minutes to one hour, was recorded and transcribed after the interviews under consent. Given the stigma surrounding EDs (Puhl and Suh, 2015), we allowed the participants to freely turn off their cameras during the interview. In the interviews, participants shared their experiences regarding how various digital device usages are associated with and result in disordered eating. We also inquired about any apps and content they perceived to affect their ED symptoms. For participants who did not respond to the survey, we also asked the survey questions during the interview. We did not bring up digital food content unless participants voluntarily mentioned it, which is consistent with our survey approach. Only the interviewees, not survey-only respondents, were rewarded approximately USD 23 for a one-hour interview.
3.3. Analysis
We followed the inductive thematic analysis (Braun and Clarke, 2012). The first author initiated the data familiarization process by transcribing the interviews and used the inductive approach to open-code the survey responses and interview transcripts. To comprehensively understand and validate the initial code, every author reviewed the data and emerging themes throughout regular discussions. Then, we developed two themes that capture digital food content types associated with their ED symptoms. Finally, we titled each theme and gave a description. Note that we report only themes that involve digital food content as other types of digital content, such as body images on social media, have been disclosed from previous studies that investigated triggering multimedia content for ED patients (Feuston et al., 2020; Gak et al., 2022).
3.4. Results
Sixteen out of twenty-three participants voluntarily connected digital food content to their disordered eating practices, revealing a nuanced relationship. Our analysis discerned two primary types of digital food content associated with disordered eating: food media and food delivery apps.
Eleven participants stated Food Media, including recipe blog posts, food pictures, food crunch or chewing ASMR videos [^2] An Autonomous Sensory Meridian Response (ASMR) refers to a tingling sensation in response to specific visual, auditory, or tactile stimuli. ASMR videos often feature soothing voices, tapping, eating, and other soothing sounds., and eating broadcasts, as associated with digital food content, reaffirming the connections between food-related content and disordered eating (Bodenlos and Wormuth, 2013; Spence et al., 2019; Choi et al., 2024). Participants’ narratives unveiled diverse motivations and practices related to food media consumption. For instance, several participants revealed they search for food media when hungry instead of actually eating, and viewing such food media triggers ED symptoms. P4 described, “I watch eating broadcasts when I am famished. For about three out of ten times, I end up binge eating. I think it is because I am overwhelmed by my hunger.” On the other hand, some participants indicated that recommended food media by multimedia platforms also stimulates their ED symptoms. P7 said, “When I repeatedly view such things (suggested food media), I unconsciously feel the urge to eat it. It would be okay if I could eat it without feeling guilty about it, but if I, even once, feel guilty (because it would lead to weight gain), it piles up on and on, pushing me to binge eat. So, in the long run, it does have an impact.
Nine participants referred to Food Delivery Apps, unveiling varied motivations and practices. Some used the apps for surfing restaurants and food pictures or ordering food to binge eat. P10 remarked, “I just wanted to look at pictures to ease my hunger. So, I used to open the [food delivery] apps.” In addition, every participant noted that “the main route to binge eating is through the food delivery apps” (P6). Most emphasized that the convenience of food delivery apps has become a pathway to binge eating. P8 said, “It’s very accessible. When I have the urge to binge eat, I don’t want to move, so I often order food through delivery apps on purpose.
Our exploratory study confirmed that many participants relate digital food content to their ED symptoms. They interacted with various food media, exhibiting multifaceted relationships ranging from alleviating hunger to triggering disordered eating. Similarly, motivations for using food delivery apps varied. These nuanced experiences, characterized by diverse motivations and consumption patterns, underscore the importance of understanding the complex relationship between digital food content and ED symptoms. This complexity motivated the subsequent interviews to further comprehend its underlying dynamics.
4. In-depth Study
4.1. Participants
We recruited 22 participants (aged 18-41, mean=26.7 years; 21 females and 1 male) through advertisement posts on the online social support communities for people with EDs under permission of the moderators in South Korea  (cop., 2018; KakaoTalk, 2022). We present participants’ demographics and ED diagnosis information in our Supplementary Materials. Two in-depth study participants, P9 and P11 (P6 and P8 in the exploratory study, respectively), also participated in our exploratory interviews. To be eligible for the study, participants were required to (1) be over 18 years old and (2) have an ED. Since many people do not seek treatment for their EDs, eligibility for this study was not contingent upon a diagnosis. Participants, however, had to identify themselves as having ED symptoms (e.g., restriction of food intake, binge eating, or compensatory behaviors). Again, by recruiting users who did not seek treatment or had discontinued it, we could shed light on the needs of users whose struggles are largely ignored in a clinical setting.
4.2. Study Protocol
All phases of our study were IRB-approved and conducted remotely due to the high social stigma of EDs (Puhl and Suh, 2015). Participants responded to a survey via email before the interview. The survey included the EDE-Q and questions about their demographics and ED symptoms. [^3] The average EDE-Q score of the participants was 4.40±1.09, compared with averages of individuals diagnosed with ED (4.02±1.28) (Aardoom et al., 2012). Following this, we conducted semi-structured interviews with all participants on Zoom, each lasting approximately thirty minutes to one hour. We allowed participants to turn off their cameras. Every interview was recorded and transcribed with consent by the participants. Our interview delved into motivations and practices of digital food content consumption by asking questions about participants’ experiences. We encouraged participants to share their varied experiences with different types of food media, covering various food videos, pictures, and posts. The detailed interview protocol is in our Supplementary Materials.
4.3. Analysis
We conducted the inductive thematic analysis (Braun and Clarke, 2012) on the responses to the descriptive questions in the survey and interview to investigate how individuals with EDs interact with digital food content and how such content relates to disordered eating behaviors. After transcribing the interviews, the first and second authors individually looked through all transcriptions to fully understand the interview content. They independently generated initial codes and, through iterative discussions on emerging themes, addressed inconsistencies, resolved disagreements, and ensured that the identified themes were strongly supported by the data. To finalize the analysis, all authors participated in comprehensive discussions. Once the analysis was completed, each theme was titled and given a description.
5. In-Depth Study Results
5.1. Patterns of Digital Food Content Consumption
Figure 1. Patterns of food media consumption.
[…]
Figure 1 depicts patterns of food media consumption. The detailed explanation of these patterns is described in Section 5.1.
Figure 2. Patterns of food delivery app use.
[…]
Figure 2 depicts patterns of food delivery app use. The detailed explanation of these patterns is described in Section 5.1.
Our participants discussed their consumption of digital food content concerning ED symptoms and how it has changed over time. Based on their experiences, recurring patterns of consuming food media, such as food images and videos, emerged, as shown in Figure 1. Initially, they started consuming food media with various motivations or simply by chance. When they first began consuming food media, they found these media satisfying their motivations and believed they had no adverse effects. However, most began to perceive that viewing food media often exacerbated their ED symptoms over time. Consequently, many would take deliberate measures to avoid watching such content. Nevertheless, there were instances where they would lapse and resume watching these media, either due to a loss of self-control or incidental circumstances.
We also identified recurring patterns of utilizing food delivery apps, as depicted in Figure 2. Initially, they engaged with these apps for various motivations, including window shopping and ordering food for planned binge eating. Remarkably, some were influenced by food media, such as eating broadcasts which often feature hosts consuming delivery food (Kang et al., 2020), which led them to use food delivery apps. As they became aware of the negative impact of using food delivery apps on their ED symptoms, many would make a conscious effort to discontinue their use (e.g., uninstall the apps). However, there were instances where they would recommence using these apps and fall back into the cycle of food delivery app usage, particularly during periods of weakened self-control.
We present their journey through identified themes followed by relevant quotations that illustrate (1) their motivations for digital food content consumption practices (Section 5.2), (2) how digital food content leads to or mitigates disordered eating (Section 5.3), and (3) the efforts made by individuals with ED to avoid digital food content and their subsequent relapses into recurring digital food content consumption patterns (Section 5.4).
5.2. Motivations for Digital Food Content Consumption
  Digital Food Content Motivation Stated Subcategories of Digital Food Content
Food Media M1 Craving for Food
eating broadcast, tasting video,
new food products showcasing video/post,
binge eating broadcast
M2 Expectations for Vicarious Satisfaction
eating broadcast, tasting video,
new food products showcasing
M3 Expectations for Comfort from Shared Experience binge eating broadcast/post
M4 Motivation for Change binge eating broadcast
M5 Expectations for Relief by Following the Media cooking video
Food Delivery Apps M6 Habitual Window Shopping
M7 Food Media-Triggered Use
M8 Orders for Disordered Eating
Table 1. Motivations for digital food content consumption. The stated subcategories of digital food content are identified from open-ended interview responses. Note that these subcategories are not always distinctly separate from one another. Detailed descriptions of each subcategory are in our Supplementary Materials.
Our interview results reaffirm the observations made in our exploratory study and deepen our understanding of motivations for consuming digital food content. We uncovered motivations for digital food content consumption practices. Table 1 outlines the themes about motivations that emerged from the in-depth study.
We identified five motivations for seeking food media. Many participants mentioned M1craving for food (N=13). P18 said, “I search for and watch food media when craving food.” The motivation of craving for food often accompanies M2expectations for vicarious satisfaction (N=7). In most cases, individuals with ED prefer to fill their hunger by watching eating scenes rather than eating food. This is similar to the general population’s motivation for watching eating broadcasts, where the focus is on deriving vicarious satisfaction from the visual and auditory experience (Anjani et al., 2020).
Some searched for videos filmed by people with EDs, such as binge eating broadcasts, with M3expectation for comfort from shared experiences (N=4). Through those videos, they sought “relief from shared experiences and feel solidarity” (P12). Every participant who remarked on expectations for relief from shared experiences noted that such media calm anxiety and stigma about having an ED. While previous research showed that the general population often watches eating broadcasts for a sense of social connectedness and attachment to the hosts (Anjani et al., 2020), our finding reveals a different dimension: individuals with EDs are sometimes drawn to videos where the host engages in disordered eating. For them, this content offers not just a connection but a sense of comfort and validation in their shared struggles. This highlights a distinct psychological motivation closely tied to their experiences with their condition. M4Motivation for change (N=2) was another motivator for watching such videos. To quote P2, “[Looking at people with EDs binge eat food] made me realize that such behavior looks very bad to other people. This made me reflect on myself. … Hence, I seek them to motivate myself to overcome my eating disorder.
Additionally, a few stated M5expectations for relief by following the media (N=2). For instance, they followed cooking videos to “alleviate the craving for food and feel relief by following the videos” (P8). Interestingly, unlike the general population who often seek practical information such as recipes or nutritional details from cooking videos (Wang et al., 2022), some participants with EDs engage with this content for emotional relief and cravings management, highlighting a unique psychological motivation.
For the use of food delivery apps, we identified three motivations. Notably, more than half of the participants who associated food delivery apps with their ED behaviors mentioned M6habitual window shopping (N=6) as a motivation for using food delivery apps. They “browse food delivery apps when bored, but do not order anything” (P4). P16 said, “it [food delivery app] is right there [on my phone], so I unconsciously open the app.
It is also notable that more than half of them mentioned M7food-media triggered use (N=6). P6 described, “After watching a video of somebody finishing a half gallon of mint chocolate ice cream, I ordered it (on a food delivery app). Often, I end up ordering and eating the same thing after watching it [eating broadcasts]. This suggests that combining different types of digital food content, specifically food media and food delivery apps, can together influence immediate behaviors among individuals with EDs. While food media on social media alone has been recognized as a powerful tool to encourage food product purchases (Ventura et al., 2021), its integration with food delivery apps appears to facilitate this effect. This dynamic interplay underscores the complex role of digital food content in shaping user behaviors, such as their food choices and eating practices.
Some participants said they use food delivery apps only to make M8orders for disordered eating (N=4), but never for normal meals. They even said they often make multiple orders for disordered eating to get enough food to binge eat and purge. P7 said, “When I binge eat, though I’m very full, I make multiple orders”.
To summarize, participants engaged in various food media, primarily expecting to quell the urge to (binge) eat to overcome EDs. On the other hand, their use of food delivery apps exhibited a range of motivations, none of which were geared toward overcoming EDs.
5.3. Digital Food Content ED-Promoting or Recovery-Empowering
     Digital Food Content How Being Associated with Disordered Eating
Food Media ED-Promoting Impact ED1 Stimulate Visually and Auditorily
ED2 Cannot Provide Vicarious Satisfaction
ED3 Present Abnormally Large Amount of Food Consumption
ED4 Trigger Overwhelming Thoughts about Food
ED5 Normalize Eating Disorders from Watching Other Patients
Recovery-Empowering Impact REC1 Curb Urge to Binge Eating by Following Recipes
REC2 Provide Relief through Shared Experiences
REC3 Promote Empowerment for Change
Food Delivery Apps ED-Promoting Impact ED1 Stimulate Visually
ED6 Lower Barrier of Getting Food
Table 2. Impact of digital food content on the urge to disordered eating: factors promoting ED and empowering recovery.
Through the interviews, we gained insights into how digital food content triggers disordered eating, mostly binge eating. We identified five themes about how food media contributes to eating disorder symptoms. Many participants discussed ED1 visual and auditory stimuli (N=11) incite their urge to binge eat. P10 said, “I can’t refrain myself from binge eating when I watch an eating broadcast with visual or auditory stimulation. The more my cravings are aroused [by watching it], the more and faster I eat.” P2 commented “the enticing display of food makes [the food media] more stimulating.” Some participants mentioned that the eating sound makes digital food media more invigorating. P17 said, “it stimulates my appetite more when YouTubers eat food with vivid sounds.
Our participants also noted that, despite their expectations for vicarious satisfaction, watching food media ED2 cannot provide them with vicarious satisfaction (N=6) but instead increases the urge for food. P19 said, “Looking back, I was falsely assuming that: oh, since the person in the video took my place and ate the food for me, I will be satisfied, which wasn’t the case, and my urges piled up, leading me to binge eat. Afterward, I regretted my actions and vomited.” Although a few studies claim that the general population gets vicarious satisfaction through watching eating broadcast (Kircaburun et al., 2021b; Anjani et al., 2020), our findings indicate that it does not always hold for people with EDs.
Notably, every participant mentioned that the ED3 abnormally large amount of food (N=22) in food media has a negative impact, such as making them normalize binge eating. P6 said, “People who eat a lot of food make me think, ‘It is okay to eat this much because there are other people who eat that much.’ ” This finding extends previous research by showing that, while large food portions in eating broadcasts can lead the general population to normalize overconsumption (Kircaburun et al., 2021c), they can also make individuals with EDs normalize their disordered eating behaviors, such as binge eating.
Among all the participants, there was a unanimous assessment of eating broadcasts, indicating at least one adverse impact. However, participants had varying opinions on the effects of other types of food media. For example, views on eating segments within non-food videos, such as vlogs, diverged. Some participants believed that eating segments were different from eating broadcasts and were not related to their ED symptoms. In contrast, others considered eating segments in non-food videos and eating broadcasts to be similar, attributing negative impacts to both.
Conflicting assessments also emerged regarding cooking posts and videos. Two participants who identified seeking relief as their motivation for watching digital food media (Section 5.2), expressed a positive view of cooking posts and videos. They found these resources helpful in REC1 curbing the urge to binge eat by following the recipes (N=2) provided. On the other hand, some other participants experienced the opposite effect, as exposure to such media led them to “continuously ruminate on the recipes and food” (P4), ED4 triggering overwhelming thoughts about food (N=5).
For binge eating broadcasts, even the same individual experienced different impacts depending on the situation or circumstances at that time. They primarily fulfilled our participants’ motivations, such as seeking REC2 relief through shared experiences (N=4) and seeking REC3 empowerment for change (N=2). Moreover, they have made promises not to binge eat after watching binge eating broadcasts. P1 said, “In the video, there are many instances where people regret and blame themselves for binge eating, so when I see that, I think, ‘I should eat in moderation, not like that.’ ” In the meantime, P6 stated that she sometimes felt that she does “not have to treat eating disorder, as many people do.” This ED5 normalization of EDs from watching other patients (N=2) could impede the recovery process for individuals with EDs.
Regarding how food delivery app use contributes to eating disorders, we identified two themes about how it assists disordered eating. One common impact between food media and food delivery apps was ED1 stimulating visually (N=2). P10 noted “With habitual use of food delivery apps, the enticing display of food stimulates my cravings.” In addition, many participants mentioned the ED6 lower barrier of getting food (N=11) on delivery apps compared with buying them in-person as a significant factor leading to disordered eating. P11 said, “Food delivery apps enabled us to conveniently obtain food. People bring food to my door only with a few clicks. So I don’t need to be self-conscious of how I look to other people. I can finish the (binge eating) process as if nothing happened. This is really bad.
In conclusion, while our participants with EDs often anticipate positive effects in overcoming EDs when they consume food media, in reality, these media frequently worsen the disorder instead. Moreover, the impact of food media varies depending on the individuals, circumstances, and its subcategories, leading to both positive and negative effects. On the other hand, when it comes to engaging with food delivery apps, our findings primarily highlight the triggering impacts associated with these apps.
5.4. Efforts to Avoid Digital Food Content Leading to Disordered Eating and Relapse
  Digital Food Content Avoidance Practices
Food Media A1Influence the Algorithm to Reduce Suggestions
Food Delivery Apps A2Uninstall Food Delivery Apps
A3Delete Accounts
A4Use Shared Accounts
Table 3. Practices for avoiding digital food content leading to disordered eating.
Eleven participants said they tried to avoid food media but often relapsed. Most mentioned that they tried to A1influence the algorithm to reduce suggestions (N=8) of food videos by unsubscribing from eating broadcast channels and/or deliberately searching for non-food content. They utilized the ‘Not interested’ and ‘Don’t recommend channel’ buttons on YouTube and the ‘Not interested’ button on Instagram to remove food media from their personalized algorithm. P11 said, “There are different measures you can take to make the (YouTube) algorithm not recommend certain videos. So, I tried to use them to remove recommendations for food-related videos.” In addition, P13 “deliberately searched for random keywords to change recommendations.
Every participant who tried to modify their suggestion algorithm noted that such efforts often fail because they get back to actively searching for food-related videos whenever they lose self-control. The algorithm quickly responds to such behaviors. P2 said, “Even after putting effort into removing food videos from my recommended feed, eventually, I give in to my urges and search for food videos. Food videos quickly showed up on my feed again.
Eleven participants reported that they tried to regulate food delivery app use. All participants mentioned A2uninstalling food delivery apps (N=11) and some stated A3deleting their accounts (N=3). However, everyone described such efforts often prove ineffective as they tend to reinstall the apps whenever they lose self-control, and reinstalling or creating new accounts is trivial. To quote P9, “I repeatedly uninstall food delivery apps trying not to binge eat, but reinstall it whenever I want to binge eat.
Two participants mentioned A4 using shared accounts (N=2) as a strategy to have an extra layer of vigilance in managing their engagement with food delivery apps. They explained that, by using shared accounts, linked with family members, they introduced an element of accountability even when they were alone. For instance, P4 shared, “If my parents are at work and I am home alone, I might order and binge eat something, and destroy the evidence, and they wouldn’t know what I did. So, there is an option for a family account. Yes, I have set that up.” This approach allowed them to create a safeguard against impulsive app usage, as their actions would be visible to others, potentially discouraging disordered eating. However, not every food delivery app offers the option of shared accounts.
In summary, our participants engaged with digital food content for various reasons. While this engagement often resulted in ED-promoting impacts, it is worth noting that some participants also used different subcategories of food media, such as cooking and binge eating broadcasts, as a way to alleviate their ED symptoms. In addition, we observed a common pattern of behavior in both food media consumption and food delivery app use, where participants tried to self-regulate their usage to prevent adverse impacts, but their efforts often went in vain.
6. Discussion
Our study highlighted the cycle of engagement and avoidance with digital food content among individuals with EDs, underscoring the importance of developing tailored support to encourage healthy interaction with digital food content for them. Herein, we share insights to break the loop of unhealthy interaction with digital food content while facilitating healthy use of digital food content (Sections 6.1 and 6.2). We also discuss future directions in digital food content curation (Section 6.3) and multimedia platform designs for vulnerable populations (Section 6.4).
6.1. Promoting Informed and Healthier Content Engagement
Figure 3. Example wireframes portraying design implications for healthier interaction with digital food content on media platforms based on our findings. 1g11g A community-based collective rating feedback, with ‘Healthy Eating’ as an example of a viewer-centric community. 1g21g A community-based content rating system where members can vote the recommendability of videos, and 1g31g A comment filtering option to highlight community-specific discussions and perspectives.
[…]
Figure 3 showcases example wireframes demonstrating community-driven feedback features on YouTube, specifically within a community named “Healthy Eating.” Circled 1 shows a community review message above a video titled, “Many Healthy Eating members rated this video not recommendable.” Circled 2 highlights a community rating feature positioned next to YouTube’s original thumbs-up and down icons, with a blue and red gradient background indicating the proportion of positive and negative votes. Circled 3 displays a “Healthy Eating” comment filter, summarizing community feedback with comments like, “Trigger viewers to have late night snack.”
Our findings underscore the detrimental impacts of digital food content on individuals with EDs. Despite efforts to avoid such content, many participants relapsed into consuming it either by chance or due to a lapse in self-control. These interaction patterns have been overlooked, leading to a lack of initiatives to promote healthy engagement with digital food content. In response, we discuss design implications of collaborative approaches to foster conscious and healthier interactions with digital food content and support individuals in maintaining self-control.
6.1.1. Community-driven Feedback on Media Platforms
Our findings reveal that individuals with EDs engage with and respond to various types of digital food content differently from the general population. While media platforms implement content warnings to protect vulnerable users (e.g., content warnings on violent content for children) (Charles et al., 2022; Center, 2024, 2021), providing warnings tailored to the diverse needs of various vulnerable populations remains a challenge (Haimson et al., 2020). This approach often overlooks less recognized harms that certain content may pose, such as the negative impacts of digital food content on individuals with EDs as identified in our study.
To address this gap, we suggest incorporating viewer community-driven feedback features into media platforms. These features could allow viewers to form communities based on shared interests or conditions, such as EDs, and enable them to see how other community members evaluate specific content (Figure 3  1g11g and  1g21g). We believe such peer-driven feedback, grounded in the concept of shared intentionality (Tenenberg et al., 2016), could provide valuable insights into potential harms not covered by standard content warnings and enhance awareness of the overlooked harms (Choi et al., 2022). By leveraging shared experiences among community members (Figure 3  1g31g), users could be better equipped to understand how particular content might affect them personally, empowering them to make more informed content choices and fostering a more supportive platform environment that addresses both well-known and emerging risks. Although platforms such as YouTube introduced community features, these are centered around content creators rather than viewers (Google, 2024; YouTube, 2024). A shift towards a more viewer-centric model could allow users to collectively rate and review content under a shared identity.
6.1.2. Supportive Food Delivery App Environment
Figure 4. Example wireframes portraying design implications for healthier interaction with digital food content on food delivery apps based on our findings. 1g11g Supportive message from a peer. 1g21g Collaborative food order with peers. 1g31g Interactive questioning aids encouraging mindfulness during the food delivery waiting period.
[…]
Figure 4 shows example wireframes illustrating features that encourage healthier use of food delivery apps. Circled 1 highlights a social feature on the app’s home tab, displaying a supportive message from Carol, ”You’re not alone in this, sweetheart. Let’s select a meal that makes you feel good inside and out,” emphasizing social support in meal selection. Circled 2 presents a social feature encouraging users to make mindful food choices through connections with acquaintances, with Carol asking, ”Hey, sweetheart. Do you mind sharing if this order is about satisfying your physical hunger, or is it more emotional?” Circled 3 depicts an intervention promoting mindful eating on the order status screen, prompting users with questions like ”What emotion am I feeling now?” and asking them to describe their emotions, fostering reflective and mindful engagement with their food choices.
We identified a recurring pattern of uninstalling and reinstalling food delivery apps among individuals with EDs to avoid unhealthy usage of these apps. To foster healthier engagement, we suggest incorporating features within the food delivery apps to promote mindful usage. For instance, a recent study demonstrated the effectiveness of reflective prompts in encouraging mindful engagement with food-related videos on YouTube for individuals with EDs (Choi et al., 2024). A similar just-in-time intervention could be applied to food delivery apps through reflective messages triggered upon app entry (Figure 4  1g11g). These prompts could encourage users to pause and reflect, potentially influencing their decisions about app use (Choi et al., 2024; Xu et al., 2022), although they have reinstalled the apps following lapses in self-control.
We also identified opportunities to enhance social features within food delivery apps to encourage healthy eating. For example, two participants used the shared account feature with family members as a way to manage their app usage, making their food-ordering through the app visible to others to avoid disordered eating. However, in the context of family informatics, it is also important to recognize ‘what is not done’ by family members (e.g., not using the app, or using it without engaging in disordered eating) (Lukoff et al., 2018). Further research could explore designs to enable family members supporting individuals with eating disorders to recognize and affirm instances of restraint or positive behavior within these apps, such as consciously opting not to use the app or using it without engaging in disordered eating. Additionally, the just-in-time prompts we suggested (Figure 4  1g11g) could be further improved by involving family members to share supportive messages (Lukoff et al., 2018; Schaefbauer et al., 2015).
Moreover, we propose enhancing food delivery apps’ shared account feature by integrating collaborative meal planning and ordering features that involve supportive relationships, such as family members. Collaborative meal planning and journaling through instant message platforms have been shown to make the decision-making process more deliberate and make people opt for healthier options, knowing that their selections are visible to others and that someone is paying attention (Kessler et al., 2015; Schaefbauer et al., 2015). By integrating this feature into food delivery apps where food choices are made, users, specifically people with EDs, could benefit from more thoughtful and socially connected meal selections, helping them make healthier utilization of food delivery apps (Figure 4  1g21g). Designers, however, should consider users’ feelings of being monitored and aim for a balanced approach between social translucence and privacy (Erickson and Kellogg, 2000).
6.2. Opportune Moments to Prevent Disordered Eating Triggered by Digital Food Content
From our study, we recognized a few moments that offer a second chance to steer users away from the impacts that promote EDs. Some participants indicated prolonged browsing or frequent engagement with food delivery apps often precedes binge eating episodes, implying opportune moments for situated intervention. These usage patterns detail the practices associated with overeating in food delivery apps (Portingale et al., 2023). We propose researchers investigate users’ varying interaction patterns with digital food content based on different motivations and impacts and design interventions to deter disordered eating behaviors.
Previous research explored digital interventions such as showing videos that teach mindful eating or promote healthy body images just before meals to encourage healthy eating behaviors (Allirot et al., 2018; Zepeda and Deal, 2008). However, these interventions often require proactive user engagement, such as manually opening and watching the videos on their devices. We propose to leverage the waiting period between ordering and food delivery on food delivery apps to provide such interventions more seamlessly.
For instance, the app could prompt users with short, guided questions that promote mindful eating, encouraging them to reflect on their feelings or hunger levels (Figure 4  1g31g). We believe these strategies could help disrupt planned binge eating, reduce impulsivity, and foster self-reflection on food ordering. Moreover, once the food arrives, the app could prompt a final reflection, such as allowing users to type a personal commitment or intention to reinforce mindful eating.
6.3. Mindful Curation of Digital Food Content for Dietary Well-being
In the field of human-food technology interaction, there has been a growing emergence of research focused on promoting healthy eating (Khot et al., 2022; Covaci et al., 2023). Simultaneously, a substantial body of research has been dedicated to improving the eating experience through enhancing the realism or stimuli of virtual food (Velasco et al., 2016; Anjani et al., 2020). Recent studies have also investigated how tech-mediated auditory stimuli, such as different music playing or manipulating eating sounds (e.g., amplifying and delaying) (Peng-Li et al., 2021; Kleinberger et al., 2023), affect users’ eating practices. Understanding the impact of various stimuli related to food, as delivered through various technologies, on eating practices is essential. Coupled with our study, we believe this understanding could contribute to creating a more health-conscious and enjoyable experience in food multimedia design and curation, ultimately enhancing users’ experiences.
Our findings stress the importance of carefully curating food multimedia content that is increasingly captivating, engaging, and often exacerbating disordered eating (Kinkel-Ram et al., 2022). Mindful digital food content curation is particularly important when promoting healthy eating among populations highly susceptible to the influential nature of digital content, such as individuals with EDs (Choi et al., 2024). Contrary to the prevailing content recommendation strategy, which reinforces users’ exposure to similar content by relying on their recent viewing history to attract users (Chaney et al., 2018; Fortune, 2019), a more mindful curation approach is needed.
For instance, multimedia platform designers could consider adopting a counterbalancing approach to prevent users from becoming trapped in the ‘filter bubble,’ where they are repeatedly exposed to the same type of content (Nguyen et al., 2014). For example, when users have heavily engaged with food media, the platform could strategically introduce content on different topics. This intentional diversification of content could help mitigate the adverse effects of triggering food media for individuals with EDs by reducing their passive exposure to food-related media. Recent research also demonstrated that reducing exposure to food-related videos on YouTube can alleviate their negative impacts on individuals with EDs, leading to fewer obsessive thoughts about foods and overall improvement in quality of life (Choi et al., 2024).
However, implementing such a counterbalancing strategy requires carefully considering the overall user experience. If not thoughtfully designed, this approach could disrupt the flow of content the users find engaging, leading to frustration or reduced platform satisfaction (Wei and Yan, 2023). Therefore, future work should explore how to balance content curation that breaks the filter bubble, supporting users’ well-being while maintaining the recommended content attractive and engaging. Additionally, future research could explore strategies for providing a counterbalance to users and assessing its impact on both user engagement and health outcomes.
6.4. Towards User-Empowered and Supportive Multimedia Platform for Vulnerable Populations
Our participants frequently described situations where, even after indicating a preference not to be recommended food content, the algorithm would revert to suggesting food-related content if they accidentally engaged once. This pattern often left users feeling frustrated and discouraged, as the platform’s algorithms seemed to negate their efforts to curtail exposure to food content. In this regard, platforms could offer users more control to tailor their content experience in alignment with their individual goals or sensitivities (Lukoff et al., 2023). We provide several design implications that empower users to curate and personalize their content exposure, fostering a sense of agency and customization that aligns with their needs and preferences.
For example, a feature that ensures the algorithm respects each user’s desired boundaries could support users in maintaining self-control and avoiding unwanted content (Baughan et al., 2022). One practical implementation could be allowing users to specify keywords that they don’t want to see in recommended content titles as keyword-based filtering has proven effective while remaining simple to implement (Choi et al., 2024). Even if a user accidentally engages with related content, this feature would prevent the algorithm from passively recommending similar content in the future.
In addition, platforms could allow users to customize, such as resetting or restoring, their content recommendation algorithms to better align with their goals and well-being (Feng et al., 2024). Our participants noted that a single failure of self-regulation nullifies all previous efforts as the algorithm again recommends food content (Shin and Lee, 2021; Nguyen et al., 2014). For instance, implementing an undo last engagement feature could help users quickly correct accidental interactions, preventing the algorithm from adjusting based on a single misstep. This approach could support ongoing control and responsiveness, strengthening users’ sense of agency in managing their content interactions on the platform.
Incorporating these designs would give users greater control over their content experience and reinforce the platforms’ commitment to promoting users’ well-being. By acknowledging the disparities arising from consuming food media and offering users tools to alleviate these challenges, platforms can create a more inclusive and empowering environment that supports users’ efforts to maintain a healthier relationship with digital food content and overcome their negative impacts.
6.5. Limitations
Our study predominantly included participants residing in South Korea, with only two individuals from North America and Europe. While food media and food delivery apps have achieved global popularity (Anjani et al., 2020; Insights, 2023), the limited demographic diversity of our participants means that our findings may not be fully generalizable across different cultural and geographic contexts.
7. Conclusion
We examined how individuals with eating disorders perceive and interact with digital food content. Our study offers insights into the complex dynamics between individuals with eating disorders and digital food content. The disparity between anticipated positive effects and actual outcomes of consuming such content underscores the need for health-conscious food content design to prevent unintended negative impacts on actual eating. Our design implications and proposed directions for mindful content curation and enhanced user control over digital food content provide a strategic foundation for HCI designers to create more tailored and supportive digital food content platforms. By addressing individuals with eating disorders’ unique needs, HCI researchers and practitioners could contribute to developing healthier and empowering digital environments. Our research contributes to the HCI discourse by examining the potential for more responsible and sensitive design practices in the realm of digital food content.
Appendix A Supplementary Material
A.1. Exploratory Study Participants Demographics
     P Age Gender
ED
Population
Diagnosis
(formal/self)
ED
Duration
EDE-Q
Score
Interview Survey Stated Digital Food Content
1 18 Female BN formal 1 years 2 months 4.83 Yes
2 19 Female BED self 4 years - Yes
3 20 Female BN formal 5 years 5.39 Yes food content, food delivery apps
4 20 Female BED self 7 years 5.48 Yes Yes food content
5 21 Female BN self 1 year - Yes
6 23 Female BN formal 4 years 4.68 Yes Yes food delivery apps
7 23 Female BN self 4 years 4.22 Yes Yes food content
8 23 Female BN formal 6 years 5 months 4.98 Yes Yes food delivery apps
9 24 Not to disclose BN self 1 year 4.32 Yes food delivery apps
10 24 Female BN formal 4 years 4.55 Yes Yes food content, food delivery apps
11 24 Female BED formal 5 years 3.97 Yes Yes food content, food delivery apps
12 24 Female BED self 9 years - Yes food delivery apps
13 25 Female BN formal 2 years 2 months 5.03 Yes food content
14 25 Non-binary BN self 3 years 4.16 Yes food content
15 26 Female BN self 3 years 3.52 Yes Yes food content, food delivery apps
16 27 Female BN formal 13 years - Yes
17 29 Male BED self 3 years - Yes food content
18 30 Female BED self 2 months - Yes food content
19 30 Not to disclose BN formal 4 years 3.98 Yes Yes food delivery apps
20 30 Female BN formal 15 years - Yes
21 36 Female BN formal 17 years 4.66 Yes
22 38 Female BED self 6 years 11 months - Yes
23 43 Female BED self 20 years - Yes food content
Table 4. Exploratory study participants demographic, eating disorder information, and stated digital food content. The listed items (i.e., food content and food delivery apps) in the stated digital food content are identified from open-ended responses in surveys and interviews. BED refers to Binge Eating Disorder and BN stands for Bulimia Nervosa. BED is characterized by recurrent episodes of excessive eating in a short timeframe (e.g., 2 hours) and feeling a loss of control over their eating behavior (Dingemans et al., 2002). BN involves a cycle of binge eating and compensatory behaviors, such as self-induced vomiting (Fairburn and Garner, 1986).
A.2. In-Depth Study Participants Demographics
     P Age Gender
ED
Population
Diagnosis
(formal/self)
ED
Duration
EDE-Q
Score
Stated Digital Food Content
1 18 Female BN self 1 year 4.86 food media, food delivery apps
2 20 Female BN self 3 months 3.84 food media
3 20 Female BN self 1 year 4.67 food media, food delivery apps
4 21 Female BN formal 2 years 4.38 food media
5 22 Female BED self 3 year 4.41 food media
6 22 Female BN formal 3 years 1 month 5.30 food media
7 22 Female BN formal 10 years 5.11 food media
8 23 Female BN self 3 years 2.72 food media, food delivery apps
9 23 Female BN formal 4 years 4.59 food media, food delivery apps
10 24 Female BED self 2 years 4.19 food media
11 24 Female BN formal 6 year 5 months 5.27 food media
12 25 Female BN self 5 years 4.82 food media, food delivery apps
13 26 Female BN self 2 years 4.71 food media
14 26 Female BED self 2 years 5.36 food media, food delivery apps
15 29 Female BN self 1 year 4.96 food media
16 29 Female BED self 2 years 4.47 food media, food delivery apps
17 33 Female BN formal 7 years 2 months 5.21 food media, food delivery apps
18 33 Female BN self 15 years 5.03 food media, food delivery apps
19 34 Female BED formal 10 years 5.03 food media, food delivery apps
20 34 Female BN formal 15 years 4.32 food delivery apps
21 39 Male BED self 10 years 2.74 food media
22 41 Female BN self 20 years 0.72 food delivery apps
Table 5. In-depth study participants demographic, ED information, and stated digital food content. The stated digital food content items are identified from open-ended interview responses. BED refers to Binge Eating Disorder and BN stands for Bulimia Nervosa. BED is characterized by recurrent episodes of excessive eating in a short timeframe (e.g., 2 hours) and feeling a loss of control over their eating behavior (Dingemans et al., 2002). BN involves a cycle of binge eating and compensatory behaviors, such as self-induced vomiting (Fairburn and Garner, 1986).
A.3. In-Depth Study Semi-structured Interview Protocol - Digital Food Content Consumption and ED Symptoms
This is an English-translated summary of the key questions we aimed to address. In the semi-structured interviews, we delved deeper and investigated topics beyond these specific questions.
Warm Up
Hello, and thank you once again for participating in this user study. As mentioned in the email, today’s interview is scheduled to last up to an hour. You are welcome to turn your camera on or off as you prefer. If there are any questions during the interview that you do not want to answer, please feel free to decline. With your consent, we will now begin recording the interview.
Exposures to digital food content and eating disorder symptoms
First, I will ask you some questions about the relationship between digital food content and the eating disorder symptoms you are experiencing.
  1. (1)
    Do you think that your digital food content consumption is associated with your ED symptoms? Why do you think it is (not) related?
  2. (2)
    What kinds of food content do you think are related?
    For each content,
    1. (a)
      What motivates you to consume the content?
    2. (b)
      Could you share any specific instances/experiences where your eating disorder symptoms led to the consumption of the content?
    3. (c)
      Are there any other factors, besides eating disorder symptoms, that drive you to consume the content?
    4. (d)
      Could you share any specific instances/experiences where consuming the content impacted your eating disorder symptoms?
    5. (e)
      Have you tried regulating your consumption of the content? If so, why and how?
Digital supports to regulate digital food content consumption
  1. (1)
    How effectively do you think digital support or interventions could prevent the negative impact or enhance the positive impact of consuming digital food content on your eating disorder symptoms? Please share any ideas or strategies you have in mind for addressing different types of digital food content.
A.4. Table of Participants-Stated Subcategories of Digital Food Content
  Subcategories of Digital Food Content Description Quote
eating broadcast (Mukbang) videos featuring the consumption of large amounts of food Eating broadcast shot by professional eating broadcast YouTubers affects me significantly.” (P1)
tasting video/post videos/posts featuring food reviews on media platforms or maps (e.g., restaurant review on Google map) When I watch food reviews, it often triggers a desire to eat, leading to urge to binge eat.” (P17)
new food products showcasing video/post videos/posts introducing new food products Watching videos where hosts taste new products makes me crave those foods, making me try them.” (P21)
binge eating broadcast/post videos/posts created by people with EDs to log eating disorder symptoms The (binge eating) blogs I’ve subscribed to is mainly about recording what they’ve eaten.” (P3)
cooking video/post videos/posts featuring the process of cooking or recipe When following cooking videos, the process from cooking to eating takes it at a more leisurely pace.” (P18)